OpenStudy (anonymous):

Solve the equation. 2x^2-5x=3

OpenStudy (j-cross18):

-3x=3 x=-1 is your answer

OpenStudy (j-cross18):

I think :P

OpenStudy (anonymous):

it can't be just substitute if you substitute -1 then the answer is 7 but we are trying to get 3

OpenStudy (anonymous):

It's a rational expression. You don't just plug the numbers in.

OpenStudy (j-cross18):

Ok umm but the answer still has to be a negative right???

OpenStudy (anonymous):

I'm not sure. It will be two numbers. I think it is 3,5

OpenStudy (anonymous):

I think it is just 3

OpenStudy (anonymous):

It can't be just 3. lol

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

are there answer choices?

OpenStudy (anonymous):

\[2x^2-5x=3\rightarrow 2x^2-5x-3=0\]Then you use the quadratic formula.\[-(-5)\pm \frac{ \sqrt{(-5)^2-4(2)(-3)} }{ 2(2) }\]Can you solve from there?

OpenStudy (anonymous):

I am confused because I haven't been working them out this way. Not quite sure where to start?

OpenStudy (anonymous):

Oops, everything should be over 2(2), including the -(-5) part

OpenStudy (anonymous):

How have you been working them out? Have you used the quadratic formula for these at all?

OpenStudy (anonymous):

I am confused too, I haven't learned any of that

OpenStudy (anonymous):

Yes I have been using the quadratic formula but I somehow was finding the right answer using a different method. I have no clue what I was doing. Whatever I was doing isn't working on this.

OpenStudy (anonymous):

Were you factoring things out, and setting them to zero? For example x^2+9x+18 -->(x+3)(x+6) then x+3=0 --> x=-3 and x+6=0-->x=-6 ? Sometimes you can't factor the quadratic out so you need to use the quadratic formula to get your zeros.

OpenStudy (anonymous):

That is what I have been doing. This can't be factored though?

OpenStudy (anonymous):

No this one can't be factored. In this case you would need to utilize the quadratic formula.

OpenStudy (anonymous):

Ok so with the formula you worked out above where would I start? -5=-25+12/4

OpenStudy (anonymous):

Alright so we have your quadratic in standard form: \[ax^2+bx+c=0\] \[2x^2-5x-3=0\] We plug your corresponding numbers into our quadratic formula: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\rightarrow \frac{ -(-5)\pm \sqrt{(-5)^2-4(2)(-3)} }{ 2(2) }\]

OpenStudy (anonymous):

We split the equation into a plus and minus version to get \[\frac{ 5+\sqrt{49} }{ 4 } and \frac{ 5-\sqrt{49} }{ 4 }\]

OpenStudy (shamil98):

2x^2-5x - 3 =0 a = 2 b = - 5 c =- 3 Using the magical quadratic formula. \[\large x = \frac{ 5 \pm \sqrt{25 - 4(2)(-3)} }{ 4 } \] \[\large x = \frac{ 5 \pm \sqrt{49} }{ 4 } = \frac{ 5 \pm 7 }{ 2 }\] \[\large x = \frac{ 5 + 7 }{ 2 } = 6\] \[\large x =\frac{ 5-7 }{ 2 } = -1\]

OpenStudy (anonymous):

So we have that \[x=\frac{ 5+7 }{ 4 }=3 ; x=\frac{ 5-7 }{ 4 }=-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Ok. That is very confusing! Thank you for your help!!

OpenStudy (anonymous):

Shamil98, you forgot to continue your 4 in your denominator, you flipped it back to 2.